OpenStudy (anonymous):

Solve for n 6n + 3 = 2?

OpenStudy (anonymous):

subtract 3 from both sides 6n=2-3 =-1 divide by 6 on both the sides n= 2-3/6

OpenStudy (anonymous):

ok thank you :{D

OpenStudy (anonymous):

be sure to best response people so their grades will go up :]

OpenStudy (anonymous):

lol ;)

OpenStudy (anonymous):

it will go up till 55-58 ..after that it crawls lol

OpenStudy (anonymous):

-1/6